OpenStudy (anonymous):

Can someone show me how to do this I'm a little confused....please, and thank you :) Solve the following system of equations using the substitution method and show all of your work. 5x + 3y = 8 y = −2x + 1

OpenStudy (jack1):

so you have y = something so sub y = something into the first equation

OpenStudy (jack1):

5x + 3y = 8 5x + 3(-2x + 1) = 8

OpenStudy (anonymous):

okay so then I distribute ?

OpenStudy (anonymous):

5x-6x+3=8.....is that wrong?

OpenStudy (jack1):

so use that to work out what x = and from that work out what y =

OpenStudy (anonymous):